OpenStudy (anonymous):

Kristen lives directly east of the park. The football field is directly south of the park. The library sits on the line formed between Kristen’s home and the football field at the exact point where an altitude to the right triangle formed by her home, the park, and the football field could be drawn. The library is 2 miles from her home. The football field is 8 miles from the library.

OpenStudy (anonymous):

A. 4 miles; \[4\sqrt{5} miles\] B. \[\sqrt{10} miles;\]\[3\sqrt{2} miles\] C. 5 miles; 9 miles D. 4 miles; \[3\sqrt{2}miles\]